what is the equation of the line that has a slope of 3 and passes through the point (9,-11)?
Answer by Alan3354(21605)

I just did one like that.
-------------
Do it like this one:
what is the equation of the line that has a slope of 6 and passes through the point (5,-12)?
-------------
Use y = mx + b to find b
-12 = 6*5 + b
b = -42
---------
y = 6x -42